OpenStudy (anonymous):

Help with this L'hospital rule problem: Solve for the following limit: Lim as x approaches 0(+) of 2sinxlnx I don't know what to do, I'm always stuck with a ln0 which is undefined no?

OpenStudy (anonymous):

@zepdrix if you don't mind helping me with this one xD

zepdrix (zepdrix):

\[\huge \lim_{x \rightarrow 0^+} 2\sin x \cdot \ln x\]As x approaches 0 from the right, the limit is approaching,\[\huge 2\cdot 0\cdot -\infty\] This is of indeterminate form. But it's not one of the forms we're allowed to apply L'Hop to. Remember we need one of these two forms,\[\huge \frac{0}{0}, \qquad \frac{\infty}{\infty}\]

zepdrix (zepdrix):

We'll have to apply some fancy trig work.\[\large \sin x=\frac{1}{\csc x}\]Substituting gives us,\[\huge \lim_{x \rightarrow 0^+} 2\sin x \cdot \ln x\quad=\quad \huge \lim_{x \rightarrow 0^+} \frac{2\cdot \ln x}{\frac{1}{\csc x}}\]

zepdrix (zepdrix):

Woops, I wrote that incorrectly :) one sec.

zepdrix (zepdrix):

\[\huge \lim_{x \rightarrow 0^+} 2\sin x \cdot \ln x\quad=\quad \huge \lim_{x \rightarrow 0^+} \frac{2\cdot \ln x}{\csc x}\]

OpenStudy (anonymous):

thats infinity / 0 . we're still not allowed to apply that?

OpenStudy (anonymous):

I mean L'hospital rule*

zepdrix (zepdrix):

Is it? Hmm lemme think.

OpenStudy (anonymous):

oh wait no thats infinity

zepdrix (zepdrix):

csc x = 1/sin x, as sin x gets closer to 0, the fraction gets bigger and bigger, Yah I think you're right, infinity.

OpenStudy (anonymous):

That isn't the answer ahaha. I hate cal 1 :(

zepdrix (zepdrix):

That's not the right path to take? :o

zepdrix (zepdrix):

Seems like we're on the right track :)

OpenStudy (anonymous):

I don't know I just have the question, no answer, I dwelled on it for a good hour nothing came to mind. I just know its not infinity :P

OpenStudy (anonymous):

* I can verify the answer but it doesn't give it per say

zepdrix (zepdrix):

No we're not done yet :) lol

zepdrix (zepdrix):

We have successfully gotten it into the form infty / infty, from here we're allowed to apply L'Hop, yes?

OpenStudy (anonymous):

yup so I got 2 1/x / csctanx which is 1/sinx * sinx/cos x which gives me 2(1/x)/(1/cosx) am i right?

OpenStudy (anonymous):

err i mean -csctanx

zepdrix (zepdrix):

-cscx cotx I think

OpenStudy (anonymous):

yup -.- so I meant 2(1/x)/-cscxcotx

zepdrix (zepdrix):

Which... after you move things around... I think we get,\[\huge \lim_{x \rightarrow 0^+}\frac{-2 \cdot \sin^2x}{x \cos x}\]

OpenStudy (anonymous):

yup and replace sin with 1-cos(square)x sorry french keyboard :P

zepdrix (zepdrix):

Hmm it looks like it might still be giving us trouble :( Hmmmm

OpenStudy (anonymous):

second derivative?

OpenStudy (anonymous):

Its now in a 0/0 form which means that you can apply L'Hop again

zepdrix (zepdrix):

But won't it just get worse..? :d

OpenStudy (phi):

you could use lim x->0 sin(x)/x = 1

zepdrix (zepdrix):

Oh derp :o good call.

OpenStudy (phi):

leaving -2 tan(x) = 0

OpenStudy (anonymous):

and phil is right xD it is so hard to see through all that. I wish i could best response both :P

zepdrix (zepdrix):

\[\large \lim_{x \rightarrow 0^+}\frac{-2 \cdot \sin^2x}{x \cos x} \quad = \quad -2\lim_{x \rightarrow 0^+}\frac{\sin x}{x} \cdot \lim_{x \rightarrow 0^+} \frac{\sin x}{\cos x}\]

OpenStudy (anonymous):

yeah I didn't see that :/

zepdrix (zepdrix):

Yah that's neato :O that limit slipped my mind heh

OpenStudy (anonymous):

Ah, thats clever